A firm sells its product in a perfectly competitive market where other firms charge a price of $80 per unit. The firm’s total costs are 202Q+Q220+2Q+Q2. The profit-maximizing output for your firm is:
Answer #1

Price = $80

Total cost = 20 + 2Q + Q2

Marginal cost = dTC / dQ , (differentiation of total cost with respect to quantity)

MC = 2 + 2Q

In perfect competition, Marginal revenue = Price = Average revenue.

The profit maximizing condition of perfect competition is at the point where Marginal revenue = Marginal cost

MR = MC

80 = 2 + 2Q

78 = 2Q

Q = 39

The profit maximizing output is 39 units.